You are here: Foswiki>FLAO Web>TowerLab>LabTest20110311 (11 Mar 2011, JavierArgomedo)Edit Attach

Tests after bringing the W2 back to the Lab

NOTE1: Flowerpot did not recognize input characters (output after power up seemed fine), touching the board and connector (flat) recovered the communication.
NOTE2: The Lamp intensity control seems to behave erratically.

There is a stain at at 45 deg clockwise that rotates with the pupil rotator 1:1

ccd39_aftercoldtest_11-03-2011.png

12:20

OFF: Cube, Lamp, Stages diabled
(plugged into lower 110V plug next to the entrance door, system turned on 30 min ago, cooling not active)
(ron test discarding noisy channel)

Bin: 1 - Framerate 990 Hz (2500 kpix/sec) - RON (e-): 8.26 (estimated)
Bin: 2 - Framerate 990 Hz ( 800 kpix/sec) - RON (e-): 6.26 (estimated)
Bin: 2 - Framerate 625 Hz ( 380 kpix/sec) - RON (e-): 4.61 (estimated)
Bin: 2 - Framerate 400 Hz ( 380 kpix/sec) - RON (e-): 4.68 (estimated)
Bin: 3 - Framerate 200 Hz ( 380 kpix/sec) - RON (e-): 4.89 (estimated)
Bin: 4 - Framerate 300 Hz ( 335 kpix/sec) - RON (e-): 4.89 (estimated)
Bin: 4 - Framerate 200 Hz ( 335 kpix/sec) - RON (e-): 4.93 (estimated)
Bin: 4 - Framerate 105 Hz ( 335 kpix/sec) - RON (e-): 5.21 (estimated)


OFF: Cube, Lamp, Stages diabled, FW, ADC, Rot, Boxfan
(ron test discarding noisy channel)

Bin: 1 - Framerate 990 Hz (2500 kpix/sec) - RON (e-): 8.20 (estimated)
Bin: 2 - Framerate 990 Hz ( 800 kpix/sec) - RON (e-): 6.15 (estimated)
Bin: 2 - Framerate 625 Hz ( 380 kpix/sec) - RON (e-): 3.93 (estimated)
Bin: 2 - Framerate 400 Hz ( 380 kpix/sec) - RON (e-): 3.98 (estimated)
Bin: 3 - Framerate 200 Hz ( 380 kpix/sec) - RON (e-): 4.04 (estimated)
Bin: 4 - Framerate 300 Hz ( 335 kpix/sec) - RON (e-): 4.00 (estimated)
Bin: 4 - Framerate 200 Hz ( 335 kpix/sec) - RON (e-): 4.12 (estimated)
Bin: 4 - Framerate 105 Hz ( 335 kpix/sec) - RON (e-): 4.19 (estimated)


OFF: Cube, Lamp, Stages diabled, FW, ADC, Boxfan
(ron test discarding noisy channel)

Bin: 1 - Framerate 990 Hz (2500 kpix/sec) - RON (e-): 8.11 (estimated)
Bin: 2 - Framerate 990 Hz ( 800 kpix/sec) - RON (e-): 6.20 (estimated)
Bin: 2 - Framerate 625 Hz ( 380 kpix/sec) - RON (e-): 4.53 (estimated)
Bin: 2 - Framerate 400 Hz ( 380 kpix/sec) - RON (e-): 4.53 (estimated)
Bin: 3 - Framerate 200 Hz ( 380 kpix/sec) - RON (e-): 4.66 (estimated)
Bin: 4 - Framerate 300 Hz ( 335 kpix/sec) - RON (e-): 4.71 (estimated)
Bin: 4 - Framerate 200 Hz ( 335 kpix/sec) - RON (e-): 4.70 (estimated)
Bin: 4 - Framerate 105 Hz ( 335 kpix/sec) - RON (e-): 4.73 (estimated)


Camera Lens chan X cable exchanged by Mauro Sozzi.

Note1: Picctrl stopped once
Note2: Bayside stage X failed to do the homing twice, had to disable/renable and retry (limit switch?).

Camera lens motor X works now!

ccd39_after_camlensconnfix_11-03-2011.png

Functional test done (Board setup etc).

-- JavierArgomedo - 11 Mar 2011
Topic revision: r2 - 11 Mar 2011, JavierArgomedo
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback